A sample of a substance containing only C and H burns in excess O2 to yield 4.4g of CO2 and 2.7g of H2O. What is the empirical formular of the compound?

1
Expert's answer
2022-02-02T21:04:02-0500

CxHy + (x + y/4)O2 = XCO2 + (y/2)H2O

4.4 2.7

according to the equation 44x = 9y

Carbon Hydrogen

mole given ratio = (44)/(44), (2.7)/(9) = 0.1,0.3

Empirical formula = (0.1)/(0.1), (0.3)/(o.1) = 1,3 ; CH3